Let A and B be two sets. (a) Show that Ac = (Ac ∩ B) ∪ (Ac ∩ Bc ), Bc = (A ∩ Bc ) ∪ (Ac ∩ Bc ).

(b) Show that (A ∩ B) c = (Ac ∩ B) ∪ (Ac ∩ Bc ) ∪ (A ∩ Bc ).

(c) Consider rolling a fair six-sided die. Let A be the set of outcomes where the roll is an odd number. Let B be the set of outcomes where the roll is less than 4. Calculate the sets on both sides of the equality in part (b), and verify that the equality holds.
Answer #1

(a) The distributive property is X cap (Y cup Z) = (X cap Y) cup (X cap Z) .

Hence, we have

(A^c cap B) cup (A^c cap B^c) = A^c cap (B cup B^c)

or (A^c cap B) cup (A^c cap B^c) = A^c cap U (for U = phi^c be the universal set)

or (A^c cap B) cup (A^c cap B^c) = A^c .

Again, we have

(A cap B^c) cup (A^c cap B^c) = (A cup A^c) cap B^c (the distributive property)

or (A cap B^c) cup (A^c cap B^c) = U cap B^c

or (A cap B^c) cup (A^c cap B^c) = B^c .

(b) We have, (A^c cap B^c) = (A cup B)^c by the De Morgan's law.

Also, we have (A^c cap B) cup (A^c cap B^c) = A^c , and hence,

(A^c cap B) cup (A^c cap B^c) cup (A cap B^c) = A^c cup (A cap B^c)

or (A^c cap B) cup (A^c cap B^c) cup (A cap B^c) = (A^c cup A) cap (A^c cup B^c)

or (A^c cap B) cup (A^c cap B^c) cup (A cap B^c) = U cap (A^c cup B^c)

or (A^c cap B) cup (A^c cap B^c) cup (A cap B^c) = (A^c cup B^c)

or (A^c cap B) cup (A^c cap B^c) cup (A cap B^c) = (A cap B)^c (De Morgan's law).

(c) We have, A= {1.3.5} and B= {1.2.3} .

(A cap B)^c = (left { 1,3,5 ight } cap left { 1,2,3 ight })^c

or (A cap B)^c = (left { 1,3 ight })^c

or (A cap B)^c = left { 2,4,5,6 ight } .

Also,

(A^c cap B) = (left { 1,3,5 ight }^c cap left { 1,2,3 ight }) or (A^c cap B) = (left { 2,4,6 ight } cap left { 1,2,3 ight }) or (A^c cap B) = left { 2 ight } ,

(A^c cap B^c) = (left { 1,3,5 ight }^c cap left { 1,2,3 ight }^c) or (AcnB) = ({2.4.6}n(4.5.6) or (A^c cap B^c) = left { 4,6 ight } ,

and (A cap B^c) = (left { 1,3,5 ight } cap left { 1,2,3 ight }^c) or (A cap B^c) = (left { 1,3,5 ight } cap left { 4,5,6 ight }) or (A cap B^c) = left { 5 ight } .

Hence, we have (A^c cap B) cup (A^c cap B^c) cup (A cap B^c) = left { 2 ight } cup left { 4,6 ight } cup left { 5 ight } or (A^c cap B) cup (A^c cap B^c) cup (A cap B^c) = left { 2,4,5,6 ight } .

Hence, we may see that (A^c cap B) cup (A^c cap B^c) cup (A cap B^c) is indeed equal to (A cap B)^c .